Science and technology. Fungicides account for 1/10 of the pesticides used in the United States. Insecticides account for 1/4 of all the pesticides used in the United States. The ratio of funigicides to insecticides used in the United States can be written 1/10 divided by 1/4. Write this ratio in simplest form.
Please help me.
Thank you.

Answer by stanbon(75887)

[1/10]/[1/4]= 4/10=2/5
Cheers,
Stan H.